A Complete Bibliography of Publications in The Computer Journal: 1950–1959

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

30 October 2025 Version 2.03

Title word cross-reference

1959 [Ano58a, Ano58d].

405 [Liv59].

access [Win59a]. Aircraft [Moo59]. Algorithms [Cle59]. America [GG59]. AMOS [Ber59]. analogue [Vin59]. Analysis [Bec58, Moo59]. Analytical [Wen59]. Application [Vin59, Wri59, HK59, HB58a, HB58b]. approach [Gea59]. art [Dou59b, Gol59]. Australia [Har59a]. Authorities [Mal59]. autocode [Bro58a, Bro58b, CF59]. Automatic [Fai58, Gil59, Har59a, Mui58, Tho58a]. avoid [Cam58].

b [Dou59b]. Binary [Gil58a]. Britain [Gol59]. British [Ano58a, Ano58c, Dou59b]. Business [Bar58, DG58, Gea58, Wri59, Gea59].

Calculation [Wil58c, Wil58b]. Carlo [Mus59]. certain [Wil59]. chemical [Gal59]. Class [Wen59]. Codiagonal [Wil58c]. coding [Ber59]. coefficients [O’B59c]. combination [Vin59]. Commerce [Dou58a, Dou58b, Dou58c, Dou59a]. commercial [Cut59, Gol59]. Complex [Str59]. Computer [Ano58a, Ano58c, Bar58, Bec58, Dou59a, Dou59c, Loo59, Moo59, Wil58a, Win59b, Wri59, The59, BB59, Bro58b, Bry59, Cut59, FdVR58, Gil58c, Gil58d, HK59, HB58a, HB58b, Win59a, YR58]. Computers [Ano58b, Dou58c, Dou58a, Dou58b, Gal59, Gol59, GG59, Mar58, Tak59, Vin59, Bro58a, Dou59b]. Computing [Har59a]. Conference [Ano58a, Ano58c, Ano58d, Har59a]. congruence [Tho58b]. conjectures [Cut59]. constant [O’B59c]. continuous [BM59]. control [Bry59, Dou58c, Dou59a, HK59]. Correspondence [BH59, Fai59, JD59, O’B59a, Ven58]. Current [Gil59].

Data [Har59a, Mal59, Dou59c, Hun59]. data-processing [Hun59]. Decade [Wil58a]. decomposition [O’B59c]. design [HB58a, HB58b]. DEUCE [BB59, Rob59]. developed [Bro58a]. Development [Tak59, Wil58a, The59]. differential [Mar58, O’B59c, Whe59]. Digital [Bec58, Moo59, Win59b, Wri59, Gil58c, Gil58d, Mar58, Tak59, Vin59].

economic [Gal59]. Eigenvectors [Wil58c, Wil58b]. electron [Vin59]. Electronic [GG59, YR58]. Elliott [Liv59]. empirical [All59]. engineering [Bry59]. equations [Mar58, O’B59c, Whe59, Wil59]. establishment [Bry59]. Evaluation [Wat59]. experience [YR58]. experiences [Gri59, Liv59]. Extraction [Gow58].

facilities [Bro58b]. factory [Bry59]. Ferranti [FdVR58, Gri59, Hun59]. films [Liv59]. First [Ano58c, Bar58, O’B59c]. Forecasting [Mui58]. Forecasts [Gea58]. Form [Gil58a]. Formula [Cle59]. Foundations [Gea58]. Four [Tho58a, YR58]. Fractions [How59]. functions [All59]. Further [Bro58b]. future [Cut59].

Generating [BM59, Tho58b]. Gray [O’B59a, O’B59b]. Gros [O’B59a, O’B59b].

Harmonic [Bec58]. HEC [Woo59]. Helicopter [Har59b]. high [Mar58]. Horner [Gil58a].

industry [Gal59]. influence [Win59a]. Information [Ano58d, Fai58]. input [FdVR58]. integrated [Cut59]. integrating [Mar58, Whe59]. Intercode [Ber59]. Interlingual [Ric58]. International [Ano58d]. interpretive [Rob59]. introduction [Bry59]. Iterative [Gow58, Wen59].

Japanese [Tak59]. June [Ano58a, Ano58d, Gol59].

Kutta [Mar58, Whe59].

Lanczos [Wil58b]. large [Cut59, Wil59]. large-scale [Cut59]. light [Bry59]. linear [O’B59c]. Loading [Sha58]. Local [Mal59].

Machine [Ric58]. magnetic [Gri59, Liv59]. management [Dou59a, GG59]. Manchester [Bro58a, Bro58b]. Mathematics [DG58]. Matrices [Wil58c, Win59b]. matrix [Wil59]. Melting [ND59]. Mercury [Bro58b, FdVR58]. Merging [JD59, Win59a]. Method [Gil58a, Gow58, Wil58b, All59, Tho58b, Whe59]. methods [Mar58, Wil59]. Model [Sha58]. modified [Tho58b]. Monte [Mus59]. multi [All59, O’B59c]. multi-order [O’B59c]. multi-variable [All59].

National [Liv59]. National-Elliott [Liv59]. next [Cut59]. Non [Wen59]. Non-Analytical [Wen59]. Note [Gow58, O’B59c, Pag58, Wat59, Whe59]. Number [Str59]. numbers [Tho58b].

off [Pag58]. Office [Tho58a]. Operations [ND59]. order [O’B59c]. ordinary [Whe59].

Parallel [Gil58b]. Paris [Ano58d]. Pegasus [CF59, Gri59, Wil59]. Perseus [Hun59]. petroleum [Gal59]. planning [Gal59, Mus59]. Plant [Ano58b]. Practice [Gil59]. Principles [Bel58, Fai59, Ven58]. problem [Mus59]. Problems [Mal59, Gil58c, Gil58d]. Process [Ano58b]. Processes [Wen59, BM59, Win59a]. Processing [Ano58d, Har59a, Mal59, Hun59]. production [Bry59, HK59, Mus59]. programmers [Gea59]. Programming [Gil58b, Gil59, How59]. programs [Bro58a, Rob59, Woo59]. pseudo [Tho58b]. pseudo-random [Tho58b].

railway [Gil58c, Gil58d]. random [Tho58b]. Rational [How59]. Recorded [Fai58]. Recording [Dou59c, Dou58c]. reduction [All59]. Reference [Dou59c]. research [YR58]. Retrieval [Fai58]. Root [Gow58, Str59]. Round [Pag58]. Round-off [Pag58]. Route [Moo59]. routine [BB59, FdVR58]. Runge [Mar58, Whe59].

Sales [Mui58]. scale [Cut59]. scheme [Ber59]. Second [Wil58a]. selection [Gea59]. separation [BM59]. Series [Wat59]. sets [Wil59]. Shop [ND59, Sha58]. simplified [Ber59]. Simulation [Har59b, ND59, Mus59]. Society [Ano58a, Ano58c]. Solution [Wil59, Gil58c, Gil58d]. Some [Har59b, Cut59]. Sorting [Bel58, Fai59, JD59, Ven58]. speed [Mar58]. Square [Str59]. state [Dou59b, Gol59]. States [GG59]. Statistical [Gea58]. statistics [YR58]. stock [Dou58c]. storage [Win59a]. strategies [BM59]. Studies [Har59b]. study [HK59]. Symposium [Gea59, Gri59, Liv59]. system [Bry59, Hun59].

Taking [Str59]. tape [Gri59, Liv59]. tapes [Gri59]. Technique [How59]. Techniques [Dou59c]. Ten [The59]. Test [Woo59]. them [Cam58]. Theory [Gil59]. time [Win59a]. tools [GG59]. tracing [Vin59]. training [Gea59]. trajectory [Vin59]. Translation [Cle59, Ric58, BB59]. Transposing [Win59b]. Trigonometric [Wat59]. tunnel [HB58a, HB58b].

United [GG59]. universities [Dou59b]. University [Bro58a]. use [Gal59, Gri59, Liv59, YR58]. user [Gea59]. Using [Bec58, Wil59].

variable [All59].

Weekly [Sha58]. Where [Cut59]. wind [HB58a, HB58b]. Work [Tho58a, Cut59].

X [Loo59]. X-1 [Loo59].

Year [Bar58, YR58]. Years [Tho58a, The59].

References

Allen:1959:MRE

[All59] C. D. Allen. A method for the reduction of empirical multi-variable functions. The Computer Journal, 1(4):196–200, January 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

       http://comjnl.oxfordjournals.org/content/1/4/196.full.pdf+
       html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
       01/Issue_04/010196.sgm.abs.html;   http://www3.oup.co.uk/
       computer_journal/hdb/Volume_01/Issue_04/tiff/196.tif;   http:
       //www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
       tiff/197.tif;   http://www3.oup.co.uk/computer_journal/hdb/
       Volume_01/Issue_04/tiff/198.tif;   http://www3.oup.co.uk/
       computer_journal/hdb/Volume_01/Issue_04/tiff/199.tif;   http:
       //www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
       tiff/200.tif.

Anonymous:1958:BCS

[Ano58a] Anonymous. British Computer Society Conference: June 1959. The Computer Journal, 1(2):70, ???? 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

       oxfordjournals.org/content/1/2/70.full.pdf+html.

Anonymous:1958:CPP

[Ano58b] Anonymous. Computers and process plant. The Computer Journal, 1(1):41, ???? 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/1/1/41.full.pdf+html.

Anonymous:1958:FBC

[Ano58c] Anonymous. First British computer society conference. The Computer Journal, 1(3):141, ???? 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/141.full.pdf+html.

Anonymous:1958:ICI

[Ano58d] Anonymous. International Conference on Information Processing Paris 15–20 June 1959. The Computer Journal, 1(3):105, ???? 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/1/3/105.full.
pdf+html.

Barnard:1958:FYB

[Bar58] A. J. Barnard. The first year with a business computer. The Computer Journal, 1(1):29–36, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/1/29.full.pdf+html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/010029.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/31.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/33.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/34.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/36.tif.

Brigham:1959:TRD

[BB59] R. C. Brigham and C. G. Bell. A translation routine for the DEUCE computer. The Computer Journal, 2(2):76–84, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/76.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/020076.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/tiff/76.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/
tiff/77.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/79.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/82.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/83.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/84.tif.

Beck:1958:HAU

[Bec58] F. Beck. Harmonic analysis using a digital computer. The Computer Journal, 1(3):117, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/117.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010117.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/117.tif.

Bell:1958:PS

[Bel58] D. A. Bell. The principles of sorting. The Computer Journal, 1 (2):71–77, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/71.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010071.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/71.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/72.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/75.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/77.tif.

Berry:1959:ISC

[Ber59] F. J. Berry. Intercode, a simplified coding scheme for AMOS.

The Computer Journal, 2(2):55–58, July 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/55.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/020055.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/tiff/55.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/
tiff/56.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/57.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/58.tif.

Bauer:1959:C

[BH59] F. L. Bauer and A. S. HouseHolder. Correspondence. The Computer Journal, 2(3):102, ???? 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/3/102.full.pdf+html.

Bernard:1959:GSC

[BM59] E. E. Bernard and P. D. A. Mole. Generating strategies for continuous separation processes. The Computer Journal, 2(2): 87–89, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/2/87.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/020087.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/87.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/88.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_02/tiff/89.tif.

Brooker:1958:APD

[Bro58a] R. A. Brooker. The autocode programs developed for the Manchester University Computers. The Computer Journal, 1(1):15– 21, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/1/15.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/010015.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/16.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/19.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_01/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/21.tif.

Brooker:1958:FAF

[Bro58b] R. A. Brooker. Further autocode facilities for the Manchester (Mercury) computer. The Computer Journal, 1(3):124–127, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/1/3/124.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010124.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/124.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/126.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/127.tif.

Bryen:1959:IES

[Bry59] J. F. A. Bryen. The introduction and establishment of a system of computer production control in a light engineering factory. The Computer Journal, 2(3):115–118, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/115.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020115.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/115.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/116.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/117.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/118.tif.

Caminer:1958:HAT

[Cam58] D. T. Caminer. — and how to avoid them. The Computer Journal, 1 (1):11–14, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/1/11.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/010011a.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/12.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/14.tif.

Clarke:1959:PA

[CF59] B. Clarke and G. E. Felton. The Pegasus autocode. The Computer Journal, 1(4):192–195, January 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/192.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010192.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/192.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/193.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/194.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/195.tif.

Cleave:1959:AFT

[Cle59] J. P. Cleave. Algorithms for formula translation. The Computer Journal, 2(2):53–54, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/2/53.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/020053.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/54.tif.

Cuttle:1959:WNS

[Cut59] G. Cuttle. Where next? Some conjectures on the future of the largescale computer in integrated commercial work. The Computer Journal, 2(2):85–86, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/2/85.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/020085.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/85.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/86.tif.

Dowse:1958:MB

[DG58] R. G. Dowse and H. W. Gearing. Mathematics in business. The Computer Journal, 1(1):22–24, April 1958. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/1/22.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/010022.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/tiff/22.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/

tiff/23.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/24.tif.

Douglas:1958:CCa

[Dou58a] A. S. Douglas. Computers and commerce: 1. The Computer Journal, 1(2):69–70, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/69.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010069.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/69.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/70.tif.

Douglas:1958:CCb

[Dou58b] A. S. Douglas. Computers and commerce: 2. The Computer Journal, 1(3):132–137, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/132.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010132.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/133.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/136.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/137.tif.

Douglas:1958:CCS

[Dou58c] A. S. Douglas. Computers and commerce: 3 — stock recording and control. The Computer Journal, 1(3):137–141, October 1958. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/3/137.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010137.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/137.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/139.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/140.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/141.tif.

Douglas:1959:CCM

[Dou59a] A. S. Douglas. Computer and commerce: 4 — management and control. The Computer Journal, 1(4):168–171, January 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/168.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010168.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/168.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/169.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/170.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/171.tif.

Douglas:1959:SAB

[Dou59b] A. S. Douglas. The state of the art — (b) Computers in British universities. The Computer Journal, 2(3):100–102, October 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/100.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020100.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/100.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/102.tif.

Douglas:1959:TRR

[Dou59c] A. S. Douglas. Techniques for the recording of, and reference to data in a computer. The Computer Journal, 2(1):1–9, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/2/1/
1.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/020001.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/1.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/2.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/4.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/7.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

02/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/9.tif.

Fairthorne:1958:ARR

[Fai58] R. A. Fairthorne. Automatic retrieval of recorded information. The Computer Journal, 1(1):36–41, April 1958. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/1/36.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/010036.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/tiff/36.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/
tiff/37.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/39.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/40.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/41.tif.

Fairthorne:1959:CPS

[Fai59] R. A. Fairthorne. Correspondence: the principles of sorting. The Computer Journal, 1(4):171, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/4/171.full.pdf+html.

Fotheringham:1958:IRF

[FdVR58] J. A. Fotheringham and M. de V. Roberts. An input routine for the Ferranti Mercury computer. The Computer Journal, 1(3):128–131, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/128.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010128.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/128.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/129.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/130.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/131.tif.

Galer:1959:UCE

[Gal59] G. S. Galer. The use of computers for economic planning in the petroleum chemical industry. The Computer Journal, 2(3):145–150, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/2/
3/145.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/020145.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_03/tiff/147.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_03/tiff/150.tif.

Gearing:1958:SFB

[Gea58] H. W. Gearing. Statistical foundations for business forecasts. The Computer Journal, 1(2):59–63, July 1958. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/2/59.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_02/010059.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/tiff/59.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/
tiff/60.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_02/tiff/62.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/tiff/63.
tif.

Gearing:1959:SST

[Gea59] H. W. Gearing. Symposium on the selection and training of programmers — 1: a business user’s approach. The Computer Journal, 2(3):107–109, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/3/107.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
020107.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/107.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/108.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
03/tiff/109.tif.

Gregory:1959:ECT

[GG59] R. H. Gregory and H. W. Gearing. Electronic computers as tools for management in the United States of America: 1956. The Computer Journal, 1(4):179–191, January 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/179.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010179.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/179.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/180.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/181.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/182.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/184.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/185.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/186.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/187.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/188.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/189.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/190.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/191.tif.

Gill:1958:BFH

[Gil58a] S. Gill. A binary form of Horner’s method. The Computer Journal, 1(2):84–86, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/84.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010084.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/84.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/85.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/86.tif.

Gill:1958:PP

[Gil58b] S. Gill. Parallel programming. The Computer Journal, 1(1):2–10, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/1/1/
2.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/010002.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/10.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/2.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_01/Issue_01/tiff/3.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/4.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/tiff/7.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/9.tif.

Gilmour:1958:SRPa

[Gil58c] A. Gilmour. The solution of railway problems on a digital computer: 1. The Computer Journal, 1(1):25–28, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/1/25.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_01/010025.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/tiff/25.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_01/
tiff/26.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_01/tiff/28.tif.

Gilmour:1958:SRPb

[Gil58d] A. Gilmour. The solution of railway problems on a digital computer: 2. The Computer Journal, 1(2):78–83, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/2/78.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_02/010078.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/tiff/78.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/
tiff/79.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_02/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/tiff/82.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/83.tif.

Gill:1959:CTP

[Gil59] S. Gill. Current theory and practice of automatic programming. The Computer Journal, 2(3):110–114, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/110.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020110.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/110.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/111.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/112.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/113.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/114.tif.

Goldsmith:1959:SAC

[Gol59] J. A. Goldsmith. The state of the art — (a) Commercial computers in Britain, June 1959. The Computer Journal, 2(3):97– 99, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/3/97.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/020097.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
03/tiff/97.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/98.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/99.tif.

Gower:1958:NIM

[Gow58] J. C. Gower. A note on an iterative method for root extraction. The Computer Journal, 1(3):142–143, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/3/142.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010142.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/142.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/143.tif.

Griffiths:1959:SEU

[Gri59] G. B. Griffiths. Symposium on experiences with the use of magnetic tape — 1: magnetic tapes on a Ferranti Pegasus. The Computer Journal, 2(3):118–119, October 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/118.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020118.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/118.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/119.tif.

Harper:1959:CAC

[Har59a] C. H. D. Harper. Conference on Automatic Computing and Data Processing in Australia. The Computer Journal, 2(3):104, ???? 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/2/3/104.full.
pdf+html.

Harrison:1959:SHS

[Har59b] J. M. Harrison. Some helicopter simulation studies. The Computer Journal, 2(1):10–23, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/1/10.full.pdf+html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/020010.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_01/tiff/10.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/11.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/12.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/14.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/15.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/16.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/tiff/21.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_01/tiff/23.tif.

Hollingdale:1958:ACWa

[HB58a] S. H. Hollingdale and M. M. Barritt. An application of a computer to wind tunnel design: 1. The Computer Journal, 1(1): 42–47, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/1/42.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/010042.sgm.abs.html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/43.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_01/tiff/44.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/45.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_01/tiff/46.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_01/tiff/47.tif.

Hollingdale:1958:ACWb

[HB58b] S. H. Hollingdale and M. M. Barritt. An application of a computer to wind tunnel design: 2. The Computer Journal, 1(2): 64–68, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/64.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010064.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/64.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/65.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/66.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/67.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/68.tif.

Hemy:1959:SAC

[HK59] D. C. Hemy and W. J. Kease. The study of the application of a computer to production control. The Computer Journal, 2(1): 24–38, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/1/24.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_01/020024.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/24.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/25.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/27.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/28.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/29.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/30.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/31.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/32.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/33.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/34.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/36.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/37.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/38.tif.

Howell:1959:NPT

[How59] K. M. Howell. A new programming technique for rational fractions. The Computer Journal, 1(4):176–178, January 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/176.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010176.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/176.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/177.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/178.tif.

Hunt:1959:FPD

[Hun59] P. M. Hunt. The Ferranti Perseus data-processing system. The Computer Journal, 2(2):68–75, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/2/68.full.pdf+html;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/020068.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/69.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/70.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/72.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/73.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/75.tif.

Jones:1959:CSM

[JD59] R. W. Jones and A. S. Douglas. Correspondence: Sorting by merging. The Computer Journal, 2(2):95–96, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/95.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/95.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/96.tif.

Livesey:1959:SEU

[Liv59] P. B. Livesey. Symposium on experiences with the use of magnetic tape — 2: magnetic films on a National-Elliott 405. The Computer Journal, 2(3):120–121, October 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/120.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020120.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/120.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/121.tif.

Loopstra:1959:XC

[Loo59] B. J. Loopstra. The X-1 computer. The Computer Journal, 2(1): 39–43, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/1/39.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_01/020039.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/39.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/40.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/43.tif.

Mallinson:1959:PLA

[Mal59] C. W. Mallinson. Problems of local authorities in data processing. The Computer Journal, 2(3):105–107, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/105.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020105.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/105.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/107.tif.

Martin:1958:RKM

[Mar58] D. W. Martin. Runge–Kutta methods for integrating differential equations on high speed digital computers. The Computer Journal, 1(3):118–123, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/118.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010118.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/118.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/119.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/120.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/121.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/122.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/123.tif.

Moorhead:1959:ARA

[Moo59] W. G. Moorhead. Aircraft route analysis on a digital computer. The Computer Journal, 1(4):160–162, January 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/160.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010160.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/160.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/162.tif.

Muir:1958:ASF

[Mui58] A. Muir. Automatic sales forecasting. The Computer Journal, 1(3):113–116, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/113.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010113.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/113.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/114.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/115.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/116.tif.

Musk:1959:MCS

[Mus59] F. I. Musk. A Monte Carlo simulation of a production planning problem. The Computer Journal, 2(2):90–94, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/90.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/020090.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/tiff/90.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/
tiff/91.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/92.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/93.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/94.
tif.

Neate:1959:SMS

[ND59] R. Neate and W. J. Dacey. A simulation of melting shop operations. The Computer Journal, 2(2):59–67, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/59.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/020059.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/tiff/59.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/
tiff/60.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_02/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/62.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/63.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_02/tiff/65.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_02/tiff/66.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/tiff/67.tif.

OBeirne:1959:CGG

[O’B59a] Thos. H. O’Beirne. Correspondence: Gray or gros? The Computer Journal, 2(2):96, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_02/Issue_02/tiff/96.tif.

OBeirne:1959:GG

[O’B59b] Thos. H. O’Beirne. Gray or gros? The Computer Journal, 2(2):96, ???? 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/2/
2/96.full.pdf+html.

OBrien:1959:NDF

[O’B59c] J. P. O’Brien. Note on decomposition into first order of multi-order linear differential equations with constant coefficients. The Computer Journal, 2(3):144, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/3/144.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
020144.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/144.tif.

Page:1958:NR

[Pag58] E. S. Page. A note on round-off. The Computer Journal, 1 (1):10, April 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/1/10.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_01/010011.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
01/tiff/10.tif.

Richens:1958:IMT

[Ric58] R. H. Richens. Interlingual machine translation. The Computer Journal, 1(3):144–147, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/144.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
010144.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/144.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_03/tiff/145.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
03/tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_03/tiff/147.tif.

Robinson:1959:DIP

[Rob59] C. Robinson. DEUCE interpretive programs. The Computer Journal, 1(4):172–176, January 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/4/172.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
010172.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_01/Issue_04/tiff/172.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_04/tiff/173.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
04/tiff/174.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_04/tiff/175.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_04/tiff/176.tif.

Shackleton:1958:MWS

[Sha58] P. Shackleton. A model for weekly shop loading. The Computer Journal, 1(2):87–89, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/87.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010087.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/87.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/88.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/89.tif.

Strachey:1959:TSR

[Str59] C. Strachey. On taking the square root of a complex number. The Computer Journal, 2(2):89, July 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/2/89.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_02/020089.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/tiff/89.tif.

Takahashi:1959:DJD

[Tak59] S. Takahashi. Development of Japanese digital computers. The Computer Journal, 2(3):122–129, October 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/122.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020122.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/122.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/123.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/124.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/125.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_03/tiff/127.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/128.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/129.tif.

ofHalsbury:1959:TYC

[The59] The Rt. Hon. The Earl of Halsbury. Ten years of computer development. The Computer Journal, 1(4):153–159, January 1959. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/153.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010153.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/153.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/154.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/155.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/156.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/158.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/159.tif.

Thompson:1958:FYA

[Tho58a] T. R. Thompson. Four years of automatic office work. The Computer Journal, 1(3):106–112, October 1958. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/3/106.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010106.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/106.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/108.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/109.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/110.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/111.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/112.tif.

Thomson:1958:MCM

[Tho58b] W. E. Thomson. A modified congruence method of generating pseudorandom numbers. The Computer Journal, 1(2):83, July 1958. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/2/83.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

01/Issue_02/010083.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/tiff/83.tif.

Venn:1958:CPS

[Ven58] J. L. Venn. Correspondence: the principles of sorting. The Computer Journal, 1(3):123, October 1958. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/1/3/123.full.pdf+html.

Vine:1959:ACA

[Vin59] J. Vine. Application of a combination of analogue and digital computers to electron trajectory tracing. The Computer Journal, 2(3):134–144, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/3/134.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
020134.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/134.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/135.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
03/tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/137.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/138.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
03/tiff/139.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/140.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/141.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
03/tiff/142.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/tiff/143.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_03/tiff/144.tif.

Watt:1959:NET

[Wat59] J. M. Watt. A note on the evaluation of trigonometric series. The Computer Journal, 1(4):162, January 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/162.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010162.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/162.tif.

Wensley:1959:CNA

[Wen59] J. H. Wensley. A class of non-analytical iterative processes. The Computer Journal, 1(4):163–167, January 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/4/163.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_04/010163.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/163.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_04/tiff/165.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_04/tiff/166.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_04/
tiff/167.tif.

Wheeler:1959:NRK

[Whe59] D. J. Wheeler. Note on the Runge–Kutta method of integrating ordinary differential equations. The Computer Journal, 2(1): 23, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/1/23.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_01/020023.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/23.tif.

Wilkes:1958:SDC

[Wil58a] M. V. Wilkes. The second decade of computer development. The Computer Journal, 1(3):98–105, October 1958. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/3/98.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010098.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/100.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/102.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/103.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/105.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/98.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/99.tif.

Wilkinson:1958:CEM

[Wil58b] J. H. Wilkinson. The calculation of eigenvectors by the method of Lanczos. The Computer Journal, 1(3):148–152, October 1958. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/3/148.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_03/010148.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/148.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/149.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_03/tiff/150.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_03/tiff/151.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_03/
tiff/152.tif.

Wilkinson:1958:CEC

[Wil58c] J. H. Wilkinson. The calculation of the eigenvectors of codiagonal matrices. The Computer Journal, 1(2):90–96, July 1958. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/1/2/90.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_02/010090.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/tiff/90.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/
tiff/91.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_01/Issue_02/tiff/92.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_02/tiff/93.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_01/Issue_02/tiff/94.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
01/Issue_02/tiff/95.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_01/Issue_02/tiff/96.tif.

Wilson:1959:SCL

[Wil59] L. B. Wilson. Solution of certain large sets of equations on Pegasus using matrix methods. The Computer Journal, 2(3):130–133, October 1959. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/2/3/
130.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_03/020130.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/tiff/130.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/tiff/131.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_03/tiff/132.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/tiff/133.
tif.

Windley:1959:ISA

[Win59a] P. F. Windley. The influence of storage access time on merging processes in a computer. The Computer Journal, 2(2):49– 53, July 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/2/49.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_02/020049.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/49.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/50.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_02/tiff/51.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
02/tiff/52.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_02/tiff/53.tif.

Windley:1959:TMD

[Win59b] P. F. Windley. Transposing matrices in a digital computer. The Computer Journal, 2(1):47–48, April 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/1/47.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_01/020047.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_01/tiff/47.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_01/
tiff/48.tif.

Woolner:1959:TPH

[Woo59] A. D. Woolner. Test programs for HEC. The Computer Journal, 2 (1):44–47, April 1959. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/2/1/44.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_01/020044.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_01/tiff/45.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_
01/tiff/47.tif.

Wright:1959:BAD

[Wri59] A. G. Wright. A business application of a digital computer. The Computer Journal, 2(3):103–104, October 1959. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/3/103.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_03/020103.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_03/tiff/103.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_03/
tiff/104.tif.

Yates:1958:UEC

[YR58] F. Yates and D. H. Rees. The use of an electronic computer in research statistics: four years’ experience. The Computer Journal, 1 (2):49–58, July 1958. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic).  URL http://comjnl.oxfordjournals.
org/content/1/2/49.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_01/Issue_02/010049.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/49.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/50.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/51.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/52.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/53.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_01/Issue_02/tiff/56.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_01/Issue_02/tiff/57.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_01/Issue_
02/tiff/58.tif.